A die is thrown twice and the sum of the numbers appearing is observed to be 8. What is the conditional probability that the number 5 has appeared at least once


1
Expert's answer
2021-07-09T05:35:26-0400

It is given that a die is thrown twice.

So the total outcome "=6^2=36."

Consider "P(A)" as the probability of getting the number "5" at least once. Consider "P(B)" as the probability of getting the sum "=8."

Sample space of B "\{(2,6), (3,5), (4,4), (5,3),(2,6)\}."


"P(B)=\dfrac{5}{36}"

Consider "P(A\cap B)" as the probability of getting the number "5" at least once and the sum equal to "8."

Sample space of "(A\cap B)=\{(3,5), (5,3)\}."


"P(A\cap B)=\dfrac{2}{36}=\dfrac{1}{18}"

The conditional probability that the number 5 has appeared at least once given that the sum is "8" is


"P(A|B)=\dfrac{P(A\cap B)}{P(B)}=\dfrac{\dfrac{1}{18}}{\dfrac{5}{36}}=\dfrac{2}{5}"